Find the dimensions of a rectangular garden which encloses 48 square feet of land if the length is 3 times the width
Answer by checkley77(12569)

L=3W
LW=48
3W*W=48
3W^2=48
W^2=48/3
W^2=16
W=SQRT16
W=4 ANS. FOR THE WIDTH.
L=3*4=12 ANS. FOR THE LENGTH.
PROOF:
4*12=48
48=48